# Deep Learning for Subspace Regression

### v2

@@ G.2 Architectures and training details
DeepONet is used for standard regression, and the intrusive technique with bases extracted from branch net. As a branch net of DeepONet we select a classical convolution network with downsampling by factor of 22 along each dimension and the increase of the number of hidden features by factor of 22. Branch net is defined by the number of features after encoder $N_{e,b}$, kernel size of convolution $k_{b}$, and number of layers $N_{b}$. Trunk net is MLP which is defined by the number of hidden neurons $N_{f,t}$, number of layers $N_{t}$ and the size of basis on the output layer $N_{\phi}$. In out grid searches we used $N_{e,b}\in[4,5]$, $k_{b}\in[3,7]$, $N_{b}=4$, $N_{f,t}=N_{\phi}\in[100,200]$, $N_{t}\in[3,4]$.
@@
PCANet is defined by the number of POD basis functions used to compress feature and targets $N_{p,f}$ and $N_{p,t}$. Number of MLP layers $N_{MLP}$ and hidden units $N_{p,MLP}$. In our experiments we use $N_{p,f}\in[100,200,300,400,500]$ for elliptic equation and $N_{p,f}\in[50,80,100]$ for Burgers equation, $N_{p,t}\in[100,200,300,400]$ for elliptic equation and $N_{p,t}\in[100,150,400]$ for Burgers equation, $N_{MLP}\in[3,4,5,6]$ for elliptic equation and $N_{MLP}\in[3,5,7]$ for Burgers equation, $N_{p,MLP}\in[100,200,300,400,500]$ for elliptic and $N_{p,MLP}\in[100,300,500]$ for Burgers equation.
@@ G.3 Additional results
We provide two additional results. For elliptic equations we compare optimality of learned bases. The results are in Figure 7(a) and Figure 7(b). Interestingly, global POD leads to a better basis than both FFNO and DeepONet. For DeepONet this is expected, since trunk net does not depend on parameters of PDE. For FFNO the result is more surprising, because the basis is extracted from the last hidden layer, so it explicitly depends on parameters. Bases of DeepPOD and subspace regression are the most optimal one, still they are highly non-optimal when compared with local POD. For the Burgers equation we compare DeepPOD and two variants of subspace regression in Table 6. SubReg(10)(10) is a subspace regression trained to approximate subspace spanned by first $10$ local POD basis functions and SubReg(5)(5) was trained with 55 local POD basis functions. DeepPOD is an unsupervised method and was trained with the whole trajectory. This implies methods are sorted from left to right in the decrease of information they receive about solutions. Interestingly, the SubReg(5)(5) – method, learning the smallest subspace – performs better almost uniformly. A possible explanation is that an optimal subspace of dimension 55 leads to good enough accuracy and is easier to learn than larger subspaces.
@@ H.2 Two-grid method
By default, the batch size is $64$, training runs for $1,000$ epochs, and each processor layer has $64$ features. We repeat this search for both $L_{1}$ and $L_{2}$ losses and for predicted subspace sizes $\{10,20,30,40\}$. We then select the top-33 hyperparameter configurations per subspace size and loss (by two-grid spectral radius) and train on the dataset with $100$ grid points. Throughout the paper, we report results for the best configuration on the dataset with $100$ grid points.

### v3

@@ G.2 Architectures and training details
DeepONet is used for standard regression, and the intrusive technique with bases extracted from branch net. As a branch net of DeepONet we select a classical convolution network with downsampling by factor of $2$ along each dimension and the increase of the number of hidden features by factor of $2$. Branch net is defined by the number of features after encoder $N_{e,b}$, kernel size of convolution $k_{b}$, and number of layers $N_{b}$. Trunk net is MLP which is defined by the number of hidden neurons $N_{f,t}$, number of layers $N_{t}$ and the size of basis on the output layer $N_{\phi}$. In out grid searches we used $N_{e,b}\in[4,5]$, $k_{b}\in[3,7]$, $N_{b}=4$, $N_{f,t}=N_{\phi}\in[100,200]$, $N_{t}\in[3,4]$.
@@
PCANet is defined by the number of POD basis functions used to compress feature and targets $N_{p,f}$ and $N_{p,t}$. Number of MLP layers $N_{MLP}$ and hidden units $N_{p,MLP}$. In our experiments we use $N_{p,f}\in[100,200,300,400,500]$ for elliptic equation and $N_{p,f}\in[50,80,100]$ for Burgers, $N_{p,t}\in[100,200,300,400]$ for elliptic equation and $N_{p,t}\in[100,150,400]$ for Burgers equation, $N_{MLP}\in[3,4,5,6]$ for elliptic equation and $N_{MLP}\in[3,5,7]$ for Burgers equation, $N_{p,MLP}\in[100,200,300,400,500]$ for elliptic and $N_{p,MLP}\in[100,300,500]$ for Burgers equation.
@@ G.3 Additional results
We provide two additional results. For elliptic equations we compare optimality of learned bases. The results are in Figure 8(a) and Figure 8(b). Interestingly, global POD leads to a better basis than both FFNO and DeepONet. For DeepONet this is expected, since trunk net does not depend on parameters of PDE. For FFNO the result is more surprising, because the basis is extracted from the last hidden layer, so it explicitly depends on parameters. Bases of DeepPOD and subspace regression are the most optimal one, still they are highly non-optimal when compared with local POD.
@@
For the Burgers equation we compare DeepPOD and two variants of subspace regression in Table 11. SubReg(10)(10) is a subspace regression trained to approximate subspace spanned by first $10$ local POD basis functions and SubReg(5)(5) was trained with $5$ local POD basis functions. DeepPOD is an unsupervised method and was trained with the whole trajectory. This implies methods are sorted from left to right in the decrease of information they receive about solutions. Interestingly, the SubReg(5)(5) – method, learning the smallest subspace – performs better almost uniformly. A possible explanation is that an optimal subspace of dimension $5$ leads to good enough accuracy and is easier to learn than larger subspaces.
@@ H.2 Two-grid method
By default, the batch size is $64$, training runs for $1,000$ epochs, and each processor layer has $64$ features. We repeat this search for both $L_{1}$ and $L_{2}$ losses and for predicted subspace sizes $\{10,20,30,40\}$. We then select the top-$3$ hyperparameter configurations per subspace size and loss (by two-grid spectral radius) and train on the dataset with $100$ grid points. Throughout the paper, we report results for the best configuration on the dataset with $100$ grid points.
@@ I.3 Results
Appendix J Sensitivity to train-test split and the choice of hyperparameters
All results in the main text are reported without error bars, that are usually computed by varying random initialisation or train-test split. All our experiments involve extensive hyperparameter search, so computing statistics for distinct initialisations or train-test split would require order of magnitude more compute. Here we demonstrate for elliptic eigenproblem (Appendix F.1) and Burgers equation (Appendix G.1) that sensetivity to train-test split is relatively small and not significant to main conclusions reached by analysis of “single-run” results.
For both Burgers and elliptic equations we randomly split dataset on train and test set $5$ times and report mean metrics and standard deviation. Networks are trained for hyperparameters found by grid search. The results are available in Table 14. Variability is clearly present, but it is not pronounced enough.
Table 14: Sensitivity to train-test split for elliptic eigenproblem and Burgers equation, subspace regression trained with $L_{2}(A,B;z)$ loss function.
elliptic, $k_{1}=k_{2}$ Burgers
$N_{\text{subspace}}$ train error $\pm$ std, $\%$ test error $\pm$ std, $\%$ train error $\pm$ std, $\%$ test error $\pm$ std, $\%$
$10$ $24.37\pm 1.47$ $30.38\pm 1.34$ $25.5\pm 1.17$ $25.71\pm 1.72$
2020 $5.69\pm 0.53$ $6.46\pm 0.26$ $1.15\pm 1.2 11$$1.0\pm 1.43$
$30$ $3.34\pm 0.26$ $3.79\pm 0.35$ $7.04\pm 0.35$ $7.01\pm 0.64$
$40$ $1.84\pm 0.12$ $2.12\pm 0.11$ $4.56\pm 0.32$ $4.72\pm 0.61$
Variability to train-test split should be compared to sensitivity to the selection of hyperparameters reported in Table 15. We see that for certain cases the span of best and worse performance reaches $100\%$ which is much higher than variability to train-test split. Note that for elliptic equation grid search was performed on downsampled dataset with $32\times 32$ grid.
Table 15: Sensitivity to train-test split for elliptic eigenproblem and Burgers equation, subspace regression trained with $L_{2}(A,B;z)$ loss function.
elliptic, $k_{1}=k_{2}$ Burgers
$N_{\text{subspace}}$ train error $\min,\max]$, $\%$ test error $\min,\max]$, $\%$ train error $\min,\max]$, $\%$ test error $\min,\max]$, $\%$
$10$ $[23.81,37.8]$ $[28.67,39.05]$ $[22.01,32.64]$ $[23.79,34.62]$
2020 $[4.74,100.0]$ $[5.51,100.0]$ $[9.53,20.14]$ $[10.06,21.05]$
$30$ $[2.77,8.6]$ $[3.33,9.06]$ $[5.58,21.81]$ $[5.84,21.7]$
$40$ $[2.06,7.05]$ $[2.31,7.34]$ $[4.3,47.43]$ $[4.31,283.33]$